two angles are complementary if their sum is 90 degrees. the second angle is 6 less then the 3times the first. find the two angles
Answer by uma(370)

Let the first angle be x
Then the second angle = 3x - 6
As they are complementary, their sum is 90 degrees
==> x + 3x - 6 = 90
==> 4x = 90 + 6
==> 4x = 96
==> x = 96/4
==> x = 24
Thus the first angle = 24 deg
Second angle = 3(24) - 6
= 66 deg

good luck!!!